OpenStudy (anonymous):

simple trig identities problem! Solve the following quadratic trig equations algebraically (without graphing technology) for the domain 0o < x < 360o. Use exact solutions whenever possible. a) 6cos^2(theta ) + cos(theta ) - 1 = 0 b) 2 - 2sin^2( theta) = cos(theta )

OpenStudy (anonymous):

put cos(theta) = x in the first equation and solve for x, and then substitute it back into cos theta

OpenStudy (anonymous):

for the second one 2-2sin^theta = 2(1-sin^2theta) = 2 cos^2theta and then put cos theta = x and solve for x

OpenStudy (anonymous):

i'm sorry but i don't understand any of that...

OpenStudy (anonymous):

substitute cos theta = x in the first problem and post what you got.

OpenStudy (anonymous):

as in.. 6x^2+x-1?? like a quadratic?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so solve that quadratic equation.

OpenStudy (anonymous):

i dont remember how to when the x^2 value has a coefficient..

OpenStudy (anonymous):

Find out and solve it. If you have done it before, I am sure you can solve it.

OpenStudy (anonymous):

http://www.youtube.com/watch?v=1Pva-Iv43Nc

OpenStudy (anonymous):

well everything becomes a fraction if i factor out the 6.... that vid didnt help.

OpenStudy (anonymous):

6x2+x-1 = 6x2+3x-2x-1

OpenStudy (anonymous):

alright, so now what is the next step?

OpenStudy (anonymous):

do you know how to factor quadratic equations?

OpenStudy (anonymous):

using the qudratic frmula? yes. why did i need to break up that middle term?

OpenStudy (anonymous):

okay. there is another way to solve quadratic eqns. its called factoring. so 6x2+3x-2x-1 = 3x(2x+1)-1(2x+1) = (3x-1)(2x+1)=0. so x = 1/3 or x = -1/2.

OpenStudy (anonymous):

ok, but how does this relate to the trig question? haha

OpenStudy (anonymous):

so if x = 1/3, it means cos theta = 1/3. and x= -1/2. so cos theta = -1/2. solve foe theta within 0 and 360 degrees

OpenStudy (anonymous):

wwhat do you mean by solving for theta between 0 and 360? sorry i'm so bad at this!

OpenStudy (anonymous):

just find theta.

OpenStudy (anonymous):

dont worry about the 0 and 360.

OpenStudy (anonymous):

post the values of theta you got.

OpenStudy (anonymous):

arent the x= values the solutions?

OpenStudy (anonymous):

No, x is just a variable you used to denote cos theta. so you have to find theta